OpenStudy (anonymous):

which value of theta satisfies the equation 2cos^2theta - cos theta=0

OpenStudy (anonymous):

\[cos^2(\theta)-cos(\theta)=0\] \[cos(\theta)(cos(\theta)-1)=0\] \[cos(\theta)=0\] or \[cos(\theta)=1\]

OpenStudy (anonymous):

oooooooops forgot the two!

OpenStudy (anonymous):

ignore my post, or just modify with the two.

OpenStudy (anonymous):

is there a formula to memorize forthis?

OpenStudy (anonymous):

\[cos(\theta)(2cos(\theta)-1)=0\] \[cos(\theta)=0\] \[cos(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

oh no the first part has nothing at all to do with trig

myininaya (myininaya):

costheta(2costheta-1)=0 cost=0 => t=pi/2,3pi/2,... 2npi+pi/2 and 2npi+3p/2 for n=0,1,2,... 2cost-1=0 => cost=1/2 => t=pi/3,11pi/6,.... 2npi+pi/3 and 2npi+11pi/6 for n=0,1,2..

OpenStudy (anonymous):

like solving \[2x^2-x=0\] \[x(2x-1)=0\] \[x=0\] \[x=\frac{1}{2}\]

OpenStudy (anonymous):

oh no im confused

OpenStudy (anonymous):

the trig part comes now.

myininaya (myininaya):

oops that 11pi/6 is wrong its 5pi/3

OpenStudy (anonymous):

the first job is to solve for \[cos(\theta)\] the next step is to solve for \[\theta\]

OpenStudy (anonymous):

the first steps have nothing at all to do with trig. no formulas, no nothing.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

think of \[cos(\theta)\] as x and solve for x. you have \[2x^2-x=0\] and this is easy to solve

OpenStudy (anonymous):

we get \[x=0\] or \[x=\frac{1}{2}\] so now \[cos(\theta)=0\] or \[cos(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

the trig part comes now, seeing if we can find numbers for theta that will make this work.

OpenStudy (anonymous):

can you do the next step, that is, can you find theta?

OpenStudy (anonymous):

im not sure/

OpenStudy (anonymous):

isnt it 0?

OpenStudy (anonymous):

ok so first off we want to find where \[cos(\theta)=0\]

myininaya (myininaya):

just look between 0 and 2pi first to see where cosine is 0

OpenStudy (anonymous):

best cheat sheet. look on it to see where cosine is 0. it is not at 0, because \[cos(0)=1\]

myininaya (myininaya):

then after just add 2npi for n=0,1,2,3,... blah blah

OpenStudy (anonymous):

cos is the x correct?

OpenStudy (anonymous):

yes we already know x=0, so therefore \[cos(\theta)=0\] now we are looking for \[\theta\] that is we are looking for an angle (number) whose cosine is 0

myininaya (myininaya):

what happens at pi/2?

OpenStudy (anonymous):

60 defgrees

OpenStudy (anonymous):

the very bottom of the cheat sheet i sent has all the angles associated with the unit circle, and all the coordinates. look on the unit circle to see where the first coordinate is 0

OpenStudy (anonymous):

0,1

OpenStudy (anonymous):

90

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

if you are working in degrees it is 90

OpenStudy (anonymous):

radians \[\frac{\pi}{2}\]

myininaya (myininaya):

is there another between 0 and 2pi, that cosine will be zero?

OpenStudy (anonymous):

but on my wksht theres noone of those answers.

OpenStudy (anonymous):

yes at 270

OpenStudy (anonymous):

whatever you prefer. so that takes care of \[cos(\theta)=0\] for the moment. now we need \[cos(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

now that one is 60

OpenStudy (anonymous):

yes 90 works, and 270 also works

OpenStudy (anonymous):

or 300 degrees

OpenStudy (anonymous):

exactly

myininaya (myininaya):

or 3pi/2 now we can get cosine is 0 when we start at these angles and go around the circle again (2pi) so the answer for cos(t)=0 is t=2npi+pi/2 and t=2npi+3pi/2

OpenStudy (anonymous):

you are working in degrees so you have the answers.

myininaya (myininaya):

that whole paragraph is still talking about the first one k?

OpenStudy (anonymous):

theo nly answers available are pi over 6 0 pi/4 or pi/m3

OpenStudy (anonymous):

then they are missing \[\frac{\pi}{2}\]

OpenStudy (anonymous):

\[\frac{\pi}{3}\] works, from the cheat sheet

OpenStudy (anonymous):

if those are your possible answers you are working in radians not degrees.

OpenStudy (anonymous):

\[2 \cos^2 \theta- \cos \theta = 0\]

OpenStudy (anonymous):

i know but i like using degrees better.

OpenStudy (anonymous):

so was pi/3 the answer? vecause cos was 1/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i semi get it,thank you

OpenStudy (anonymous):

i guess it just asked 'which one of these is a solution"

OpenStudy (anonymous):

did you look at the cheat sheet? because it is a good one

OpenStudy (anonymous):

yeah which value satisfies the the equation

OpenStudy (anonymous):

yes, it had a lot of good material thank you

OpenStudy (anonymous):

could you helo me with my other math questions?

OpenStudy (l):

2 cos_th^2 - cos_th = 0 cos_th (2 cos_th - 1) = 0 -> cos_th =0 or 1/2 -> th = theta = +pi/2 or - pi/2 or -pi/3 or +pi/3 *

OpenStudy (anonymous):

@LynFran